OpenStudy (anonymous):

a roller coaster has a mass of 500kg contains 15 people each with average mass of 67 kg at lowest point on its run the roller coaster is 5.0m off the ground at the highest point it is 62 above the ground. A.) calculate the gravitational energy of the loaded roller coaster at its highest ponit. B) Assuming that the track frictionless what is speed of the roller coaster at the lowest point on the track. Plzzzzzzzzzzzzz someone help me

OpenStudy (schrodingers_cat):

The total weight of the roller coaster is 500+(15)(67)= 1505 The gravitational potential energy at 62 m is mgy or (1505)(9.8)(62) = 914438 The speed at 5 m above the ground is found by mgy=1/2mv^2 as the gravitational energy at the peak has to be converted into kinetic energy. So, v=sqrt(2gh) h= 62-5= 57 v= sqrt((2)(9.8)(57)= 33.42 m/s